# Tell my why: Training preferences-based RL with human preferences and step-level explanations

### v1

@@ Abstract
Human-in-the-loop reinforcement learning (HRL) allows the training of agents through various interfaces, even for non-expert humans. Recently, preference-based methods (PBRL), where the human has to give his preference over two trajectories, increased in popularity since they allow training in domains where more direct feedback is hard to formulate. However, the current PBRL methods have limitations and do not provide humans with an expressive interface for giving feedback. With this work, we propose a new preference-based learning method that provides humans with a more expressive interface to provide their preference over trajectories and a factual explanation (or annotation of why they have this preference). These explanations allow the human to explain what parts of the trajectory are most relevant for the preference. We allow the expression of the explanations over individual trajectory steps. We evaluate our method in various simulations using a simulated human oracle (with realistic restrictions), and our results show that our extended feedback can improve the speed of learning. Code & data: github.com/under-rewiev
@@ 1 Introduction
Our work is also motivated by the concept of "scaffolding" (van de Pol et al. (2010), which human teachers often use when teaching tasks to students. While scaffolding contains many techniques, like breaking the task into sub-parts, using verbal cues, or relating the content to other knowledge, we focus on a single technique in this work. Teachers often highlight essential features of the tasks (or solutions) to the student. This highlighting is often called factual explanations.
@@ 4 Implementation
Our main contribution is the extension of the feedback possibilities in PBRL for humans. In addition to their preferences between the two trajectories, the human can now explain which timesteps they deems essential for their decision. This evaluative judgment, accompanied by detailed binary explanations provided by the human over timesteps, is recorded the dataset $\mathcal{D}$ as a quintuple $(\sigma_{0},\sigma_{1},y,\text{e}_{1},\text{e}_{2})$. Here, $\text{e}_{1}$ and $\text{e}_{2}$ directly represent binary vectors of human-provided explanations, each having the same length as the corresponding segments $\sigma_{0}$ and $\sigma_{1}$. These binary vectors, $\text{e}_{1}$ and $\text{e}_{2}$, serve to elucidate the significance of individual timesteps within their respective segments, providing binary indications (0 or 1) of the human’s perceived importance at each step. Figure 1 shows an overview of our approach.
@@ 4.1 Reward Learning with annotated preferences
$L_{\text{annotation}}=\frac{1}{2N}\sum_{e,\sigma\in\{(e_{0},\sigma_{0})(e_{1},\sigma_{1})\}\in\mathcal{D}}[-(e\cdot\log(f(\xi(\sigma)))+(1-e\cdot\log(1-f(\xi(\sigma)))]$ (5)

### v2

# Tell me why: Training preference-based RL with human preferences and step-level explanations

@@ Abstract
Human-in-the-loop reinforcement learning allows the training of agents through various interfaces, even for non-expert humans. Recently, preference-based methods (PbRL), where the human has to give his preference over two trajectories, increased in popularity since they allow training in domains where more direct feedback is hard to formulate. However, the current PBRL methods have limitations and do not provide humans with an expressive interface for giving feedback. With this work, we propose a new preference-based learning method that provides humans with a more expressive interface to provide their preference over trajectories and a factual explanation (or annotation of why they have this preference). These explanations allow the human to explain what parts of the trajectory are most relevant for the preference. We allow the expression of the explanations over individual trajectory steps. We evaluate our method in various simulations using a simulated human oracle (with realistic restrictions), and our results show that our extended feedback can improve the speed of learning.
@@ 1 Introduction
Our work is also motivated by the concept of "scaffolding" (van de Pol et al. (2010)), which human teachers often use when teaching tasks to students. While scaffolding contains many techniques, like breaking the task into sub-parts, using verbal cues, or relating the content to other knowledge, we focus on a single technique in this work. Teachers often highlight essential features of the tasks (or solutions) to the student. This highlighting is often called factual explanations.
@@ 4 Implementation
Our main contribution is the extension of the feedback possibilities in PBRL for humans. In addition to their preferences between the two trajectories, the human can now explain which timesteps they deems essential for their decision. This evaluative judgment, accompanied by detailed binary explanations provided by the human over timesteps, is recorded the dataset $\mathcal{D}$ as a quintuple $(\sigma_{0},\sigma_{1},y,\text{e}_{0},\text{e}_{1})$. Here, $\text{e}_{0}$ and $\text{e}_{1}$ directly represent binary vectors of human-provided explanations, each having the same length as the corresponding segments $\sigma_{0}$ and $\sigma_{1}$. These binary vectors, $\text{e}_{0}$ and $\text{e}_{1}$, serve to elucidate the significance of individual timesteps within their respective segments, providing binary indications (0 or 1) of the human’s perceived importance at each step. Figure 1 shows an overview of our approach.
@@ 4.1 Reward Learning with annotated preferences
$L_{\text{annotation}}=\frac{1}{2N}\sum_{e,\sigma\in\{(e_{0},\sigma_{0}),(e_{1},\sigma_{1})\}\in\mathcal{D}}[-(e\cdot\log(f(\xi(\sigma))))+(1-e\cdot\log(1-f(\xi(\sigma))))]$ (5)